# Large-$N$ SU(4) Schwinger boson theory for coupled-dimer antiferromagnets

### v2

@@ III.2 Path-integral formulation
The path-integral formulation is implemented by introducing identities expressed in terms of the $\mathrm{SU}(4)$ coherent states of each dimer, rather than the more conventional approach using $\mathrm{SU}(2)$ coherent states of each spin [41, 45]:
@@
where $\tau$ denotes the imaginary time, $\beta$ the inverse of temperature, $\mathscr{H}(\bar{b},b)=\mathscr{H}_{0}(\bar{b},b)+\mathscr{H}^{\prime}(\bar{b},b)$ the Hamiltonian density, obtained from ^0+^subscript^$\hat{\mathscr{H}}_{0}+\hat{\mathscr{H}}^{\prime}$ by replacing the operators $\hat{b}_{j,\mu}^{\dagger},\hat{b}_{j,\mu}$ with complex fields $\bar{b}_{j,\mu},b_{j,\mu}$ that parametrize the CP3 manifold of $\mathrm{SU}(4)$ coherent states. The additional fields are the Lagrange multipliers $\lambda_{j}$, which enforce the local constraint given in Eq. (29).
@@
$V^{\alpha}_{k,k+q}={1\over 2}\sum_{a^{\prime},\delta}{\cal J}^{a^{\prime}a}_{\delta}\left((V_{\delta}^{a^{\prime}})^{\dagger}e^{-i{\bm{k}}\cdot{\bm{\delta}}}+P(V_{\delta}^{a^{\prime}})^{*}Pe^{i({\bm{k}}+{\bm{q}})\cdot{\bm{\delta}}}\right),$
where $\alpha$ denotes the component of fluctuation field $\Phi_{j,\delta}^{a}$ (see below for definition of matrix $P$), ${\bm{\delta}}$ is the bond vector of bond $\delta$,
$V^{\alpha}_{k,k+q}=-{1\over 2}\sum_{a^{\prime},\delta}{\cal J}^{aa^{\prime}}_{\delta}\left(V_{\delta}^{a^{\prime}}e^{i({\bm{k}}+{\bm{q}})\cdot{\bm{\delta}}}+P(V_{\delta}^{a^{\prime}})^{T}Pe^{-i{\bm{k}}\cdot{\bm{\delta}}}\right),$
where $\alpha$ denotes the component of fluctuation field $\bar{\Phi}_{j,\delta}^{a}$, and
$V^{\alpha}_{k,k+q}={i\over 2}I_{2N}.$ (64)
where $\alpha\equiv\lambda$. Note that, to simplify the diagrammatic calculation, we have symmetrized the internal vertex function by implementing the particle-hole symmetry of the Nambu field, $\psi_{k}=P(\psi_{-k}^{\dagger})^{T}$, with $P=(\sigma_{x}\otimes I_{N})$.
@@ IV Saddle point solution
Before delving into the SP solution for the broken-symmetry state, it is worthwhile to compare our SP solution for the QPM with the ’bond operator’ approach introduced by Sachdev and Bhatt [30]. The approximation proposed by Sachdev and Bhatt can be derived from our large-$N$ expansion if we approximate the expectation values of the ${S}_{j,\delta}$ and ${T}_{j,\delta}$ link fields in Eq. (75) in the following way:
$\langle_{j,\delta}\rangle_{\rm sp}\approx\langle b_{j,0}\rangle_{\rm sp}\langle b_{j+\delta,0}\rangle_{\rm sp},$ $\langle_{j,\delta}^{\;}\rangle_{\rm sp}\approx\langle b^{\dagger}_{j+\delta,0}\rangle_{\rm sp}\langle b_{j,0}\rangle_{\rm sp},$
@@ V Magnetic excitation spectrum
$u_{S}^{\mu}=u_{+}^{\mu}+u_{-}^{\mu},\ \ u_{A}^{\mu}=u_{+}^{\mu}-u_{-}^{\mu}.$ (125)
along with the simplified notation for the magnetic susceptibility, $\chi_{\alpha}^{\mu\nu}$, where the subscript $\alpha=S,A$ denotes the symmetric or antisymmetric channels, respectively. Due to the $\mathrm{SU}(2)$ rotation symmetry, or the residual $\mathrm{U}(1)$ symmetry about the local moments in the Néel order phase, the magnetic susceptibility is diagonal in the spin index. The diagonal components satisfy $\chi_{\alpha}^{xx}(q)=\chi_{\alpha}^{yy}(q)=\chi_{\alpha}^{zz}(q)$ in the QPM phase and $\chi_{\alpha}^{xx}(q)\neq\chi_{\alpha}^{yy}(q)=\chi_{\alpha}^{zz}(q)$ in the Néel ordered phase, assuming that the local moments align along the $x$ direction.
@@ V.1 Saddle-point approximation
where the prime symbol indicates that the sum is restricted to momenta ${\bm{k}}$ such that $\varepsilon_{{\bm{k}},n}>0,\varepsilon_{{\bm{k}}+{\bm{q}},m}>0$. This contribution gives rise to the two-particle continuum in the DSSF  [^3] In the SP approximation, the two-particle continuum corresponds to the excitation of a singlet boson and a triplet boson (see Eq. (130)). However, this is a spurious continuum, as the singlet boson is not a physical mode. The physical continuum appears at higher energies and corresponds to a three-triplon excitations, which are only obtained after properly accounting for quantum fluctuations (this a generic feature of the large-$N$ theories). In contrast, the two-particle continuum in the symmetric channel of the DSSF is physical, as it consists of two triplon excitations..
@@
Figure 9: Longitudinal DSSF in the anti-symmetric channel for $g=0.8$. (a) SP result including two $\delta$-peaks corresponding to two unphysical modes. (b) Result after including the counter-diagram shown in Fig. 11 (b). The lines indicate the single-boson dispersion obtained from the SP Hamiltonian. The dotted square denotes the region where an extra low-intensity quasi-flat mode emerges from the fluctuations of the auxiliary fields. (c) Cancellation of the $\delta$-peak of the SP susceptibility upon adding the the counter-diagram shown in Fig. 11 (b). (d) Low-frequency scaling of the intensity of the continuum at the ordering wavevector $(\pi,\pi)$. The black dashed line indicates the slope associated with $1/\omega$ scaling.
@@ V.2 Amplitude fluctuation in Néel AFM phase
where

### v3

@@ III.2 Path-integral formulation
The path-integral formulation is implemented by introducing identities expressed in terms of the SU($N=n^{2}$) coherent states of each dimer, rather than the more conventional approach using $\mathrm{SU}(n)$ coherent states of each spin [44, 45]:
@@
where $\tau$ denotes the imaginary time, $\beta$ the inverse of temperature, $\mathscr{H}(\bar{b},b)=\mathscr{H}_{0}(\bar{b},b)+\mathscr{H}^{\prime}(\bar{b},b)$ the Hamiltonian density, obtained from ^0+^subscript^$\hat{\mathscr{H}}_{0}+\hat{\mathscr{H}}^{\prime}$ by replacing the operators $\hat{b}_{j,\mu}^{\dagger},\hat{b}_{j,\mu}$ with complex fields $\bar{b}_{j,\mu},b_{j,\mu}$ that parametrize the CPN-1 manifold of $\mathrm{SU}(N=n^{2})$ coherent states. The additional fields are the Lagrange multipliers $\lambda_{j}$, which enforce the local constraint given in Eq. (27).
@@
$V^{a}_{k,k+q}={1\over 2}\sum_{a^{\prime},\delta}{\cal J}^{a^{\prime}a}_{\delta}\left((V_{\delta}^{a^{\prime}})^{\dagger}e^{-i{\bm{k}}\cdot{\bm{\delta}}}+P(V_{\delta}^{a^{\prime}})^{*}Pe^{i({\bm{k}}+{\bm{q}})\cdot{\bm{\delta}}}\right),$
where $a$ denotes the component of fluctuation field $\Phi_{j,\delta}^{a}$ (see below for definition of matrix $P$), ${\bm{\delta}}$ is the bond vector of bond $\delta$,
$V^{a}_{k,k+q}=-{1\over 2}\sum_{a^{\prime},\delta}{\cal J}^{aa^{\prime}}_{\delta}\left(V_{\delta}^{a^{\prime}}e^{i({\bm{k}}+{\bm{q}})\cdot{\bm{\delta}}}+P(V_{\delta}^{a^{\prime}})^{T}Pe^{-i{\bm{k}}\cdot{\bm{\delta}}}\right),$
where $a$ denotes the component of fluctuation field $\bar{\Phi}_{j,\delta}^{a}$, and
$V^{a=\lambda}_{k,k+q}={i\over 2}I_{2N}.$ (66)
Note that, to simplify the diagrammatic calculation, we have symmetrized the internal vertex function by implementing the particle-hole symmetry of the Nambu field, $\psi_{k}=P(\psi_{-k}^{\dagger})^{T}$, with $P=(\sigma_{x}\otimes I_{N})$.
@@ IV Saddle point solution
Before delving into the SP solution for the broken-symmetry state, it is worthwhile to compare our SP solution for the QPM with the ’bond operator’ approach introduced by Sachdev and Bhatt [30]. The approximation proposed by Sachdev and Bhatt can be derived from our large-$N$ expansion if we approximate the expectation values of the ${S}_{j,\delta}$ and ${T}_{j,\delta}$ link fields in Eq. (LABEL:eq:speqs) in the following way:
$\langle S_{j,\delta}\rangle_{\rm sp}\approx\langle b_{j,0}\rangle_{\rm sp}\langle b_{j+\delta,0}\rangle_{\rm sp},$ $\langle T_{j,\delta}^{\;}\rangle_{\rm sp}\approx\langle b^{\dagger}_{j+\delta,0}\rangle_{\rm sp}\langle b_{j,0}\rangle_{\rm sp},$
@@
For completeness, we computed the mean-field phase diagram in the $(g,\kappa)$ plane, which is shown in Fig. 7. For the physical case, $\kappa=1/4$, the phase diagram features both quantum paramagnetic (QPM) and antiferromagnetic (AFM) phases. Reducing $\kappa$ destabilizes both of these phases.
Within the QPM region, decreasing $\kappa$ reduces the condensate fraction of the singlet boson, $n_{c,0}$, eventually driving a transition into a gapped quantum spin liquid (QSL) phase once $n_{c,0}$ vanishes. In the AFM phase, a similar reduction in $\kappa$ leads to a sequential suppression of the boson condensates: the triplet condensate $n_{c,\pi}$ vanishes at the AFM-to-QPM transition, followed by the disappearance of the singlet condensate $n_{c,0}$ at the subsequent QPM-to-QSL transition.
In the QSL phase, the saddle-point solution is invariant under a staggered ${\mathrm{U}}(1)$ gauge transformation, defined by $b_{j,\mu}\to b_{j,\mu}$ for sites $j\in{\cal A}$ and $b_{j,\mu}\to-b_{j,\mu}$ for sites $j\in{\cal B}$. In accordance with this symmetry, the saddle-point expectation values of the link operators $\hat{T}_{jk}$ and $\hat{B}_{jk}$, which connect the ${\cal A}$ and ${\cal B}$ sublattices, vanish. This staggered gauge symmetry is spontaneously broken upon entering the QPM phase, as indicated by the condensation of the singlet boson $b_{j,0}$.
Figure 7: Saddle-point phase diagram on the $g-\kappa$ plane.
@@ V Magnetic excitation spectrum
$S_{j,S}^{\mu}=S_{j,+}^{\mu}+S_{j,-}^{\mu},\ \ S_{j,A}^{\mu}=S_{j,+}^{\mu}-S_{j,-}^{\mu}.$ (127)
along with the simplified notation for the magnetic susceptibility, $\chi_{\alpha}^{\mu\nu}$, where the subscript $\alpha=S,A$ denotes the symmetric or antisymmetric channels, respectively. Accordingly, we introduced the symmetric and antisymmetric external vertices, $u_{S}^{\mu}=u_{+}^{\mu}+u_{-}^{\mu}$ and $u_{A}^{\mu}=u_{+}^{\mu}-u_{-}^{\mu}.$ Due to the $\mathrm{SU}(2)$ rotation symmetry, or the residual $\mathrm{U}(1)$ symmetry about the local moments in the Néel order phase, the magnetic susceptibility is diagonal in the spin index. The diagonal components satisfy $\chi_{\alpha}^{xx}(q)=\chi_{\alpha}^{yy}(q)=\chi_{\alpha}^{zz}(q)$ in the QPM phase and $\chi_{\alpha}^{xx}(q)\neq\chi_{\alpha}^{yy}(q)=\chi_{\alpha}^{zz}(q)$ in the Néel ordered phase, assuming that the local moments align along the $x$ direction.
@@ V.1 Saddle-point approximation
where the prime symbol indicates that the sum is restricted to momenta ${\bm{k}}$ such that $\varepsilon_{{\bm{k}},n}>0,\varepsilon_{{\bm{k}}+{\bm{q}},m}>0$. This contribution gives rise to the two-particle continuum in the DSSF  [^3] In the SP approximation, the two-particle continuum corresponds to the excitation of a singlet boson and a triplet boson (see Eq. (132)). However, this is a spurious continuum, as the singlet boson is not a physical mode. The physical continuum appears at higher energies and corresponds to a three-triplon excitations, which are only obtained after properly accounting for quantum fluctuations (this is a generic feature of the large-$N$ theories). In contrast, the two-particle continuum in the symmetric channel of the DSSF is physical, as it consists of two triplon excitations..
@@
Figure 10: Longitudinal DSSF in the anti-symmetric channel for $g=0.8$. (a) SP result including two $\delta$-peaks corresponding to two unphysical modes. (b) Result after including the counter-diagram shown in Fig. 12 (b). The lines indicate the single-boson dispersion obtained from the SP Hamiltonian. The dotted square denotes the region where an extra low-intensity quasi-flat mode emerges from the fluctuations of the auxiliary fields. (c) Cancellation of the $\delta$-peak of the SP susceptibility upon adding the counter-diagram shown in Fig. 12 (b). (d) Low-frequency scaling of the intensity of the continuum at the ordering wavevector $(\pi,\pi)$. The black dashed line indicates the slope associated with $1/\omega$ scaling.
@@ V.2 Amplitude fluctuation in Néel AFM phase
where the subscript “fl” denotes “correction due to fluctuations” of the auxiliary fields,